I'm trying to find the asymtotes of a hyperbola equation. What is the square root of 6 over 2 times 3?
Answer by lwsshak3(11628)

I'm trying to find the asymtotes of a hyperbola equation. What is the square root of 6 over 2 times 3?
**
√6/(2*3)=√6/6
leave as is, cannot be further simplified.
..
asymptotes of hyperbolas are straight lines so you can used the standard form of this equation:
y=mx+b, m=slope, b=y-intercept
The slope depends on whether hyperbola has a vertical or horizontal transverse axis. Standard form of equation for a hyperbola with horizontal transverse axis: (x-h)^2/a^2-(y-k)^2/b^2=1. For vertical transverse axis: (y-k)^2/a^2-(x-h)^2/b^2=1. Slope for a horizontal transverse axis hyperbola=ħb/a and ħa/b for a vertical transverse axis hyperbola. To find b, the y-intercept, use the x,y coordinates of the center thru which the asymptotes pass.